The sum of first and second term of a geometric progression is 3 and the sum of

3rd and 4th terms is 12. Find the first term and the common ratio of the progression
and state the first four terms of the progression.
1
Expert's answer
2020-11-16T17:45:23-0500

Let the first and second term be a and r respectively.

The terms up to the 4th term will be: a, ar, ar^2, ar^3.

The sum of the first and second terms will, therefore, be a+ar=3 simplified as a(1+r)=3

The sum of the third and the fourth term will, therefore, be ar^2 + ar^3=12 simplified as ar^2 (1+r)=12

As seen, the common factor is (1+r) and this easily calls for the changing of both equations to make this the common factor. The equations will, therefore, be (1+r)=3a and (1+r)=12ar^2

We proceed to equate these equations as a simplification to find the unknowns. This will be 3a=12ar^2 that breaks down to 1/4 = r^2

Finding the square root of both sides gives the answer that the common ration r =1/2 and also r=-1/2.

If r=1/2 then,

According to the formula a+ar=3, changes to after replacing r to get a+1/2a=3, a= 2

The first four terms will be a, ar, ar^2, ar^3, which is 2, (2*1/2), (2*1/2^2), (2*1/2^3), which is 2, 1, 1/2, 1/4


The Answers are, therefore:

The First Term is 2

The Common Ratio is 1/2

The First Four Terms are 2, 1, 1/2, 1/4


On the other hand, if r =-1/2 then using the equation a(1+r)=3, then a=3/(1+-1/2), which is a=3/0.5=6

If r = -1/2 then a=6

The first four terms will therefore be, a, ar, ar^2, ar^3, which is 6, -3, 1.5 and -3/4.

If r=-1/2 the Answers are, therefore:

The First Term is 6

The Common Ratio is is r as -1/2

The First Four Terms are 6, -3, 1.5, -3/4